OpenStudy (anonymous):

6>or equal to 1/4(3+x)

OpenStudy (anonymous):

is it \[6 \ge \frac{3+x}{4}\] ?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

its a quarter

OpenStudy (anonymous):

lemme rewrite it

OpenStudy (anonymous):

what you want for answer?

OpenStudy (anonymous):

solve for x??

OpenStudy (anonymous):

\[6\ge 1/4 \left( 3+x \right)\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and yes solve for x

OpenStudy (anonymous):

bill96:its the same as i wrote!

OpenStudy (anonymous):

no the 1/4 is a fraction, the (3+x) is a separate equation multiplied by 1/4

OpenStudy (anonymous):

\[\frac{1}{4}\cdot(3+x)=\frac{3+x}{4}\]

OpenStudy (anonymous):

......no....

OpenStudy (anonymous):

ok one second

OpenStudy (anonymous):

how do you use that freakin equation button, its so damn complicated

OpenStudy (anonymous):

\[6\ge1/4 (3+x)\]

OpenStudy (anonymous):

i know, its complicated..

OpenStudy (anonymous):

ok the 3+x IS NOT IN THE DENOMINATOR

OpenStudy (anonymous):

x=(-∞,21] is this the type of answer you lookin for??

OpenStudy (anonymous):

so,u have \[6\geq\frac{3+x}{4}\implies24\geq(3+x)\implies x\leq21\]

OpenStudy (anonymous):

if its not, im probably over thinking this..

OpenStudy (anonymous):

NO NO NO omg...the 3+x is not in the fraction at ALL...

OpenStudy (anonymous):

its in parentheses after the fraction, which is 1 over 4

OpenStudy (anonymous):

bill96: u have \[\frac{1}{4}(3+x)=\frac{3+x}{4}\]

OpenStudy (anonymous):

THAT's what it is.

OpenStudy (anonymous):

ok..bill the answer is THIS: \[6\geq\frac{1}{4}(3+x)\implies24\geq(3+x)\implies x\leq21\]

OpenStudy (anonymous):

Ok, thanks.

OpenStudy (anonymous):

ya my answer was to complicated xD